If the demand function faced by a firm is:

Q = 90 – 2P

TC = 2 + 57Q – 8Q2 + Q3


Determine the level of output at which the firm maximizes the profit.


Determine the best level of output for the above question by the MR and MC approach.



1
Expert's answer
2021-04-28T09:34:26-0400

a) The firm will maximize profits at a point where MR = MC

TR = PQ = (45 - Q/2)"\times" Q

=> MR = 45 - Q

TC = 2 + 57Q - 8Q2 + Q3

=> MC = 57 - 16Q + 3Q2

MR = MC

=> 45 - Q = 57 - 16Q + 3Q2

=> 3Q2 - 15Q + 12 = 0

=> Q2 - 5Q + 4 =0

=> Q2 - Q - 4Q + 4 =0

=> Q(Q-1) -4(Q-1) = 0

=> Q = 4 or Q = 1

Thus, the answer is Q = 4

b) Q = 90 – 2P

"P=45-1\/2Q"

"MR=dP\/dQ=-1\/2"

"MC=dTC\/dQ=57-16Q+3Q^2"

Now MC=MR

"57-16Q+3Q^2=-1\/2"

"6Q^2-32Q+115=0"

On solving quadratically,

"Q_1=2.67 and \space Q_2" "=2.67"

so 2.67 is the best level of output